MaterialGrowl:

Izzy spends 6.85 on icecream and apples. The cost of the apples= 0.95$ The cost of the icecream= 0.99$ by using compatible numbers, find how many apples and icecream she bought.

Krissy525:

Is that all there is to the problem, or is there more?

MaterialGrowl:

Thats all there is Its a decimal division problem btw-

Krissy525:

@materialgrowl wrote:
Thats all there is Its a decimal division problem btw-
ohhh, so it doesnt say how many total items she bought or anything

MaterialGrowl:

Yeah

xXAikoXx:

First, what do you know about compatible numbers?

MaterialGrowl:

Absolutely nothing :D

Krissy525:

Isn't it where u estimate?

@xxaikoxx wrote:
First, what do you know about compatible numbers?

xXAikoXx:

@materialgrowl wrote:
Absolutely nothing :D
Well, you have to know the definition of compatible numbers in order for me to help.

MaterialGrowl:

Ok :D

MaterialGrowl:

In mathematics, compatible numbers are the numbers that are easy to add, subtract, multiply, or divide mentally. Compatible numbers are close in value to the actual numbers that make estimating the answer and computing problems easier.

xXAikoXx:

@materialgrowl wrote:
In mathematics, compatible numbers are the numbers that are easy to add, subtract, multiply, or divide mentally. Compatible numbers are close in value to the actual numbers that make estimating the answer and computing problems easier.
Good job, but next time put it in your own words.

MaterialGrowl:

Okk

Krissy525:

that means $6.85 would go to $6.90, and $0.99 would be $1 and $0.92 would be $0.90 right?

MaterialGrowl:

I dont think you estimate 0.99

Hananko:

add? or no?

Krissy525:

I thought you would tho

MaterialGrowl:

@hananko wrote:
add? or no?
Its division

Krissy525:

I may be wrong tho, its been so long

Hananko:

ok

Hananko:

you have to line up the decimals or it will be not right when you do the math on paper

Hananko:

*very time*

Krissy525:

its 0.95 to $1 not 0.92 to 0.90

MaterialGrowl:

mhm-

Krissy525:

Would it be $6.85 to $7 then?

Hananko:

um i think so

Krissy525:

@jhonyy9 could we pls get some help

Mercury:

Something that stands out to me about this problem: the last digit of the total cost is 5. the only way to get a digit ending in 5 from an item costing 0.99 is to have 5 of it. 5*0.99 = 4.95 from there: 6.85 - 4.95 = 1.9 coveniently, 1.9 / 2 = 0.95, or the cost of 1 apple. so $1.9 would be two apples. therefore, 5 units of ice cream + 2 apples gives us a total of $6.85.